Documenta Mathematica

Journal der Deutschen Mathematiker-Vereinigung

Band 4, 1999

Andrew Ranicki Singularities, Double Points,

Controlled Topology and Chain Duality  1–59

Detlev W. Hoffmann On a Conjecture of Izhboldin

on Similarity of Quadratic Forms  61–64

Caterina Consani, with an Appendix by Spencer Bloch The Local Monodromy

as a Generalized Algebraic Correspondence  65–108

R. Weikard On Rational and Periodic Solutions

of Stationary KdV Equations  109–126

B. Kreußler Twistor Spaces With a Pencil

of Fundamental Divisors  127–166

Bruno Kahn Les Classes de Chern Modulo p

d’une Représentation Régulière  167–178

R. Illner and S. Rjasanow Difference Scheme

for the Vlasov-Manev System  179–201

Ahmed Laghribi Sur les Formes Quadratiques de Hauteur 3

et de Degré au Plus 2  203–218

Bernadette Perrin-Riou Théorie d’Iwasawa

et Loi Explicite de Réciprocité  219–273

Marcel Griesemer, Roger T. Lewis, Heinz Siedentop A Minimax Principle for Eigenvalues in Spectral Gaps:

Dirac Operators with Coulomb Potentials  275–283

Kengo Matsumoto Presentations of Subshifts

and Their Topological Conjugacy Invariants  285–340

U. Haagerup and S. Thorbjørnsen Random Matrices and K-Theory

for Exact C^{∗}-Algebras  341–450

Marco Brunella On the Automorphism Group

of a Complex Sphere  451–462

Karsten Matthies A Subshift of Finite Type in the Takens-Bogdanov Bifurcation

with D_{3} Symmetry  463–485

M. Lübke Einstein Metrics and Stability for Flat Connections on Compact Hermitian Manifolds, and a Correspondence

with Higgs Operators in the Surface Case  487–512

Eberhard Kirchberg and Simon Wassermann

Permanence Properties of C*-exact Groups  513–558

Marc A. Rieffel

Metrics on State Spaces  559–600

Lutz Mattner

What Are Cumulants ?  601–622

J. Piontkowski and A. Van de Ven The Automorphism Group of Linear Sections

of the Grassmannians 𝔾(1, N)  623–664

Sorin Popa Some Properties of the Symmetric Enveloping Algebra of a Subfactor,

with Applications to Amenability and Property T  665–744